OpenStudy (anonymous):

find the exact value fir tan(6pi) please i need this to be explained

OpenStudy (perl):

Did you try tan(6pi ) = sin(6pi) / cos(6pi)

OpenStudy (perl):

since tan x = sin x / cos x

OpenStudy (anonymous):

where did you get that from?

OpenStudy (perl):

thats the definition of tan x

OpenStudy (anonymous):

trig functions

OpenStudy (lyrae):

 Trig definitions: \[\sin \theta = \frac{ A }{ C } \rightarrow A = C \sin \theta \]\[\cos \theta = \frac{ B }{ C } \rightarrow B = C \cos \theta\]\[\tan \theta = \frac{ A }{ B } = \frac{ C \sin \theta }{ C \cos \theta } = \frac{ \sin \theta }{ \cos \theta }\]